$SPAD/src/input complex.input

Timothy Daly December 3, 2016

Abstract

Contents

2

1 License

— license —

--Copyright The Numerical Algorithms Group Limited 1991.

———-

— * —

)set break resume
)spool complex.output
)set message test on
)set message auto off
)clear all
-- Input generated from ComplexXmpPage
--S 1 of 16
a := complex(4/3,5/2)
--R
--R
--R        4   5
--R   (1)  - + - %i
--R        3   2
--R                                             Type: Complex(Fraction(Integer))
--E 1
--S 2 of 16
b := complex(4/3,-5/2)
--R
--R
--R        4   5
--R   (2)  - - - %i
--R        3   2
--R                                             Type: Complex(Fraction(Integer))
--E 2
--S 3 of 16
a + b
--R
--R
--R        8
--R   (3)  -
--R        3
--R                                             Type: Complex(Fraction(Integer))
--E 3
--S 4 of 16
a - b

--R
--R
--R   (4)  5%i
--R                                             Type: Complex(Fraction(Integer))
--E 4
--S 5 of 16
a * b
--R
--R
--R        289
--R   (5)  ---
--R         36
--R                                             Type: Complex(Fraction(Integer))
--E 5
--S 6 of 16
a / b
--R
--R
--R          161   240
--R   (6)  - --- + --- %i
--R          289   289
--R                                             Type: Complex(Fraction(Integer))
--E 6
--S 7 of 16
% :: Fraction Complex Integer
--R
--R
--R        - 15 + 8%i
--R   (7)  ----------
--R         15 + 8%i
--R                                             Type: Fraction(Complex(Integer))
--E 7
--S 8 of 16
3.4 + 6.7 * %i
--R
--R
--R   (8)  3.4 + 6.7 %i
--R                                                         Type: Complex(Float)
--E 8
--S 9 of 16
conjugate a
--R
--R
--R        4   5
--R   (9)  - - - %i

--R        3   2
--R                                             Type: Complex(Fraction(Integer))
--E 9
--S 10 of 16
norm a
--R
--R
--R         289
--R   (10)  ---
--R          36
--R                                                      Type: Fraction(Integer)
--E 10
--S 11 of 16
real a
--R
--R
--R         4
--R   (11)  -
--R         3
--R                                                      Type: Fraction(Integer)
--E 11
--S 12 of 16
imag a
--R
--R
--R         5
--R   (12)  -
--R         2
--R                                                      Type: Fraction(Integer)
--E 12
--S 13 of 16
gcd(13 - 13*%i,31 + 27*%i)
--R
--R
--R   (13)  5 + %i
--R                                                       Type: Complex(Integer)
--E 13
--S 14 of 16
lcm(13 - 13*%i,31 + 27*%i)
--R
--R
--R   (14)  143 - 39%i
--R                                                       Type: Complex(Integer)
--E 14

--S 15 of 16
factor(13 - 13*%i)
--R
--R
--R   (15)  - (1 + %i)(2 + 3%i)(3 + 2%i)
--R                                             Type: Factored(Complex(Integer))
--E 15
--S 16 of 16
factor complex(2,0)
--R
--R
--R                      2
--R   (16)  - %i (1 + %i)
--R                                             Type: Factored(Complex(Integer))
--E 16
)spool
)lisp (bye)

———-

References^{[1] nothing}

7